OpenStudy (anonymous):

anyone good in Calculus please help Expand the "horizons" of this problem:

OpenStudy (anonymous):

\[(a). F(x)=\int\limits_{x}^{0}f(t) dt \]

OpenStudy (anonymous):

\[\int_x^0=-\int_0^x \]

OpenStudy (anonymous):

\[(b). F(x)=\int\limits_{0}^{x}\left| f(t) \right|dt\]

OpenStudy (anonymous):

@cliffsedge yes but how do i show it on the graph

OpenStudy (anonymous):

I don't really understand what your question is. These look like applications of FTC.

OpenStudy (anonymous):

\[(c). F(x)=\int\limits_{x}^{0}f(\left| t \right|)dt\]

OpenStudy (anonymous):

it is but i dont know to expand the horizons of the problems

OpenStudy (anonymous):

I'm not sure what the expression "expand the horizons" means in this context.

OpenStudy (anonymous):

i dont know it either but thanks anyways